The coupon collector problem calculates the expected number of days it takes to get n different coupons, if one receives one of the n coupons at random each day. The number of days is approximately n(0.577 + ln n). Use this to calculate the expected number of TCP connections a random port scan (scanning port numbers 0 through 1023) needs to eventually check all 1,024 well- known port numbers.
Answer #1

ANSWER:

  • substitute n in the given equation n(0.577 + in n) with n = 1024 which will be expected number of random port scan that need to cover all 1024 port numbers.
  • Hence expected number of TCP connection needed = 1024(0.577 + ln 1024) = 1024*(0.577 + 6.932) = 1024*7.508 = 7688.192
  • Hence expected number of TCP connection needed = 7689